# Retrieval-Augmented Semantic Parsing: Using Large Language Models to Improve Generalization

### v1

@@ Abstract
Open-domain semantic parsing remains a challenging task, as models often rely on heuristics and struggle to handle unseen concepts. In this paper, we investigate the potential of large language models (LLMs) for this task and introduce Retrieval-Augmented Semantic Parsing (RASP), a simple yet effective approach that integrates external lexical knowledge into the parsing process. Our experiments not only show that LLMs outperform previous encoder-decoder baselines for semantic parsing, but that RASP further enhances their ability to predict unseen concepts, nearly doubling the performance of previous models on out-of-distribution concepts. These findings highlight the promise of leveraging large language models and retrieval mechanisms for robust and open-domain semantic parsing.
Retrieval-Augmented Semantic Parsing:
Using Large Language Models to Improve Generalization
Xiao Zhang University of Groningen xiao.zhang@rug.nl                        Qianru Meng Leiden University q.r.meng@liacs.leidenuniv.nl                        Johan Bos University of Groningen johan.bos@rug.nl
@@ 1 Introduction
Open-domain semantic parsing involves mapping natural language text to formal meaning representations, capturing the concepts, relations between them, and the contects in which they appear Oepen and Lønning (2006); Hajič et al. (2012); Banarescu et al. (2013); Bos et al. (2017); Martínez Lorenzo et al. (2022). Such meaning representations are applied in many downstream applications—ranging from database querying to embodied questions answering—where parsers must handle a vast array of concepts that may not appear in the training data. While neural encoder-decoder architectures have shown impressive performance in semantic parsing tasks, their reliance on training distributions constrains their ability to generalize, especially to out-of-distribution (OOD) concepts.
Most existing semantic parsers struggle to interpret unseen words or rare senses, often defaulting to the most frequent meaning encountered during training. As a result, they fail to adapt to novel linguistic phenomena and remain limited to fixed patterns. Recent work Zhang et al. (2024a) have attempted to mitigate these limitations by encoding concept representations numerically, forcing models to learn underlying ontologies from resources like WordNet Fellbaum (1998). However, these approaches require substantial preprocessing and intricate encodings that can be difficult for models to fully exploit.
@@
  • How can these large language models be leveraged to improve generalization to out-of-distribution concepts? Beyond simple architecture comparisons, we investigate whether LLMs can be guided to handle concepts more flexibly, using their ability to interpret and integrate external information.
In Section 2 we provide background on the semantic formalism of our choice, earlier approach to semantic parsing, and the challenge of an important sub-task, word sense disambiguation. Then we propose Retrieval-Augmented Semantic Parsing (RASP) in Section 3, a technique that integrates a retrieval mechanism into the parsing process. RASP leverages external lexical knowledge (e.g., WordNet) in the input, enabling the model to dynamically access and interpret relevant concept information. By incorporating this retrieval step (Section 4), we relax the reliance on lemma-based mappings and allow the model to adapt more naturally to unseen words or senses. Our results show that this approach nearly doubles the performance on predicting OOD concepts compared to previous methods, demonstrating a substantial advancement in handling challenging open-domain data (Section 5).
@@ 2.1 Discourse Representation Structure
Figure 1: Three formats of Discourse Representation Structure (DRS) for "Mary didn’t commit a crime.": the box notation, a directed acyclic graph, and the sequence notation.
@@ 2.2 Semantic Parsing
Early approaches to semantic parsing primarily relied on rule-based systems Woods (1973); Hendrix et al. (1977); Templeton and Burger (1983). The advent of neural methodologies, coupled with the availability of large semantically annotated datasets Banarescu et al. (2013); Bos et al. (2017); Abzianidze et al. (2017), marked a significant shift in semantic parsing techniques Barzdins and Gosko (2016); van Noord and Bos (2017); Bevilacqua et al. (2021a). The introduction of pre-trained language models within the sequence-to-sequence framework further improved parsing performance van Noord et al. (2018, 2020); Ozaki et al. (2020); Samuel and Straka (2020); Shou and Lin (2021); Bevilacqua et al. (2021a); Zhou et al. (2021); Martínez Lorenzo et al. (2022); Zhang et al. (2024b); Liu (2024a, b); Yang et al. (2024). Furthermore, several studies introduced more pre-training tasks specifically designed for semantic parsing Bai et al. (2022); Wang et al. (2023a). With the rise of large language models, there has been considerable discussion about leveraging these models for semantic parsing, achieving notable results through techniques like prompting and chain-of-thought reasoning Roy et al. (2022); Ettinger et al. (2023); Jin et al. (2024). However, there is currently no work that leverages the knowledge and understanding capabilities of large language models to address the generalization problem in semantic parsing.
@@ 2.3 Word Sense Disambiguation
The generalization problem introduced in the previous section can also be understood as word sense disambiguation (WSD) for out-of-distribution concepts, within the context of semantic parsing. For instance, consider the sentence "She had £10,000 in the bank", with the target word "bank". In traditional WSD tasks, a predefined inventory of possible senses (e.g., 1. sloping land; 2. financial institution; 3. a long ridge or pile; 4. …) would be provided, and the WSD model would classify the word according to one of these senses Navigli (2009); Bevilacqua et al. (2021b).
In semantic parsing, WSD can be seen as a subtask Zhang et al. (2024a), but it is more challenging because the parsing model must generate the correct sense directly without access to an explicitly provided sense inventory. However, traditional knowledge-based WSD offers a potential solution that inspires our approach: by retrieving and presenting all possible concepts as alternatives, we can explicitly provide external information to the model, thereby enhancing its generalization capability. As a consequence, this requires the model to be able to process long contexts, making the LLMs become the preferences, in particular retrieval augmented generation.
@@ 3 Retrieval-Augmented Semantic Parsing
Gold DRS female.n.02 Name "Mary" time.n.08 TPR now birdwatch.v.01 Agent -2 Time -1 ELABORATION <1 female.n.02 ANA -3 see.v.01 Experiencer -1 Time +1 Stimulus +3 time.n.08 TPR now harrier.n.03 golden_eagle.n.01 entity.n.01 Sub -2 Sub -1 Sub +1 hobby.n.03
@@
Different from the Dense Passage Retrieval Karpukhin et al. (2020) method, which is commonly employed in question-answering tasks, our retrieval process is designed to be more straightforward and tailored to the needs of semantic parsing. The process begins with tokenizing and lemmatizing [^1] https://www.nltk.org/api/nltk.stem.wordnet.html the source text. Following these, we perform a search for relevant concept synsets in an external knowledge base, specifically WordNet. For example, in the sentence "Mary went for birdwatching. She saw a harrier, a golden eagle, and a hobby", the retrieval process would identify multiple synsets for "go", "birdwatch", "see", "harrier", "golden eagle" and "hobby", as illustrated in Table 1. Additionally, to ensure comprehensive coverage of multi-word expressions, which are critical in capturing the correct semantic meaning, we employ a hierarchical n-gram search strategy. This strategy involves sequential searches using 4-gram, 3-gram, 2-gram, and 1-gram patterns, thereby ensuring that no multi-word expressions (such as "golden eagle") are overlooked.
@@
By incorporating retrieved concepts, RASP goes beyond relying solely on the input sequence and training data, adding additional context to guide generation. For example, when handling words with multiple meanings, like "hobby," retrieved synsets help the model select the correct interpretation based on glosses and context. This integration sharpens the model’s focus on relevant concepts, reducing the likelihood of generating incorrect or overly broad outputs, particularly for out-of-distribution concepts.
@@ 4.1 Datasets
We conduct our experiments on the Parallel Meaning Bank (PMB, version 5.1.0) [^3] https://pmb.let.rug.nl/releases Abzianidze et al. (2017); Zhang et al. (2024b). We first use the gold-standard English data of the PMB to evaluate the large language models and their retrieval-augmented version under in-distribution conditions.
To further assess the models’ ability to handle out-of-distribution (OOD) concepts, we adopt the challenge set proposed by Zhang et al. (2024a), which is also derived from the PMB. Neural semantic parsers often default to the first sense of unknown concepts–an approach that can lead to "lucky guesses" without truly understanding new words. The challenge set, consisting of 500 sentences, is deliberately designed to eliminate this shortcut. Each sentence includes at least one concept that does not appear in the training data and does not correspond to the first sense in the ontology. In total, the challenge set contains 410 unknown nouns, 128 verbs, and 65 modifiers (adjectives and adverbs). By evaluating on this set, we measure the true generalization capability of the models, testing whether they can correctly interpret novel concepts rather than relying on heuristic assigning.
@@ 4.2 Experiment Settings
It is crucial to note that large language models, when used in zero-shot or few-shot scenarios, tend to perform poorly on the highly complex graph structures inherent in formal meaning representations such as DRS. Prior work Ettinger et al. (2023); Zhang et al. (2024a) demonstrates that without fine-tuning, LLMs struggle to match the performance of models specifically optimized for these tasks. Therefore, in our experiments, we fine-tune all large language models.
@@
Due to computational constraints, we select open-sourced LLMs with model sizes under 10B parameters, including phi3-4B, Mistral-7B, LLaMa3-3B, LLaMa3-8B, Gemma2-2B, Gemma2-9B, Qwen2.5-3B, and Qwen2.5-7B. These models strike a balance between state-of-the-art language understanding and manageable resource requirements. For fine-tuning, we employ Low-Rank Adaptation (Hu et al., 2021, LoRA), a parameter-efficient technique that introduces trainable low-rank matrices ($A$ and $B$) into the model’s layers, greatly reducing computational overhead:
${W+\Delta W=W+A\cdot B}$ (3)
We compare our results against several strong baselines, including BART, T5, byT5, TAX-parser Zhang et al. (2024b), and AMS-Parser Yang et al. (2024), all of which were previously fine-tuned on PMB data. We exclude work conducted on earlier versions of PMB or using silver data. Additionally, we do not apply retrieval augmentation to these baseline models due to input length constraints, which limit their ability to incorporate external lexical sources efficiently.
@@
Table 3: Performance of baseline models, large language models (Normal) and their retrieval-augmented variants (RASP) on standard test, with percentage changes in parentheses. Size is the number of model’s parameters (B: billion). IFR is Ill-Formed Rate and the number of ill-formed prediction are in parentheses. Note: AMS-Parser Yang et al. (2024) performs well for IFR for it is a compositional neuro-symbolic system. TAX-Parser Zhang et al. (2024a) is a neuro-symbolic system, trained with a novel encoded meaning representation.
@@ 4.3 Evaluation Metrics
We used SMATCH and its variants to evaluate the performance of the models. SMATCH Cai and Knight (2013), referred to as Hard-SMatch, strictly matches concepts, where any discrepancy results in a non-match. In contrast, its variant, Soft-SMatch Opitz et al. (2020), considers concept similarity when matching. Instead of adopting the approach of using word-embedding similarity, we applied the Wu-Palmer similarity Wu and Palmer (1994), as introduced by Zhang et al. (2024b). Wu-Palmer similarity provides a precise measure of semantic similarity between concepts based on their positions within the WordNet taxonomy. Unlike embedding-based methods, it does not rely on external training and easily adapts to changes in WordNet’s structure or content. The calculation is:
@@ 5.2 Performance on the Challenge Set
Among the baselines, TAX-Parser Zhang et al. (2024a) stands out, achieving an overall similarity score of 39.68. However, some Normal (non-RASP) large language models already exceed this performance on the challenge set. For example, Gemma2-9B (Normal) obtains an overall score of 42.54, indicating that LLMs can yield improvements, even without retrieval augmentation. When retrieval augmentation (RASP) is introduced, these large language models show substantial additional gains. For example, Gemma2-9B (RASP) achieves an overall similarity score of 70.41, compared to the best baseline’s 39.68—an increase of over 30 absolute points. These gains are particularly remarkable for noun concepts, with relative improvements of approximately 70% to 95%. Verbs show increases between about 25% and 77%, and modifiers improve by roughly 26% to 43%.
These results directly support our second research question regarding improving out-of-distribution generalization. While model scaling alone can yield moderate improvements, the integration of external lexical knowledge through retrieval allows LLMs to select more accurate concepts in OOD scenarios. In effect, RASP helps the models "look up" relevant information, enhancing their concept selection and producing more semantically appropriate results. In this case, retrieval-augmented LLMs not only outperform strong baselines like TAX-Parser but also set the state-of-the-art for OOD semantic parsing performance.
@@ 5.3 Error Analysis on the Challenge Set
We picked 22 instances, comprising 11 completely perfect predictions (WuP=1.00) and 11 imperfect predictions (WuP<1.00) made by RASP, as presented in Table 5. With respect to the perfect predictions, it is evident that the retrieval significantly enhance the model’s ability to interpret most out-of-distribution concepts. For instance, in the text about birdwatching, the word "hobby" clearly refers to a species of bird. The model without RAG defaults to the most frequent sense number, predicts hobby.n.01 (an auxiliary activity). In contrast, retrieval provides the glosses of each sense related to the noun "hobby" and leads the model to pick hobby.n.03 (a falcon), by explicit lexical connections between "falcon" in the gloss of hobby.n.03 and the context provided by "birdwatching".
@@
The verb "alarm" in Table 5 is an instance of the similarity problem. The challenge arises because some of its senses have similar glosses, such as alarm.v.01 (fill with apprehension or alarm) and alarm.v.02 (warn or arouse to a sense of danger). Similar issues occur with the verbs "wax", "order", "observe" and "hobble". Although glosses were carefully crafted by lexicographers, they don’t always show a clear difference in meaning Mihalcea and Moldovan (2001); Navigli (2006).
In cases of insufficient textual context, such as with the noun "kite" in the sentence "We saw a kite gliding in the sky", the sense annotators chose kite.n.04 (a bird of prey). However, kite.n.03 (a plaything) could perhaps also be appropriate given the limited context provided by this sentence. Similar issues can be raised in the sentences with the noun "tripper" and the verb "feathering".
The third cause can be attributed to the model’s linguistic coverage. A case in point is "pen": the meanings of pen.n.01 (a writing implement) and pen.n.05 (a female swan) are quite different, but the latter is the correct one in the text "Jane saw two swans. The elegant pen glided gracefully across the tranquil lake". However, the model fails to distinguish them, likely because "pen" is rarely used to refer to "swan" in available corpora. As a result, the models may not have encountered this sense during training, making it challenging for them to predict a meaning they have not been exposed to. In sum, while retrieval drastically improves concept prediction, there are still some difficulties that can pose challenges for the models.

### v2

# Retrieval-Augmented Semantic Parsing: Improving Generalization with Lexical Knowledge

@@ Abstract
Open-domain semantic parsing remains a challenging task, as neural models often rely on heuristics and struggle to handle unseen concepts. In this paper, we investigate the potential of large language models (LLMs) for this task and introduce Retrieval-Augmented Semantic Parsing (RASP), a simple yet effective approach that integrates external symbolic knowledge into the parsing process. Our experiments not only show that LLMs outperform previous encoder-decoder baselines for semantic parsing, but that RASP further enhances their ability to predict unseen concepts, nearly doubling the performance of previous models on out-of-distribution concepts. These findings highlight the promise of leveraging large language models and retrieval mechanisms for robust and open-domain semantic parsing.
@@ 1 Introduction
Open-domain semantic parsing involves mapping natural language text to formal meaning representations, capturing the concepts, relations between them, and the contexts in which they appear Oepen and Lønning (2006); Hajič et al. (2012); Banarescu et al. (2013); Bos et al. (2017); Mart´ınez Lorenzo et al. (2022). Such meaning representations are applied in many downstream applications—ranging from database querying to embodied questions answering—where parsers must handle a vast array of concepts that may not appear in the training data. While neural encoder-decoder architectures have shown impressive performance in semantic parsing tasks, their reliance on training distributions constrains their ability to generalize, especially to out-of-distribution (OOD) concepts.
Most existing semantic parsers struggle to interpret the symbols, such as rare senses, often defaulting the unseen words to the most frequent meaning encountered during training. As a result, they fail to adapt to novel linguistic phenomena and remain limited to fixed patterns. Recent work Zhang et al. (2025) have attempted to mitigate these limitations by encoding concept representations numerically, forcing models to learn underlying symbolic knowledge from resources like WordNet Fellbaum (1998). However, these approaches require substantial preprocessing and intricate encodings that can be difficult for models to fully exploit.
@@
  • How can these large language models be leveraged to improve the generation of out-of-distribution concepts? Beyond simple architecture comparisons, we investigate whether LLMs can be guided to handle concepts more flexibly, using their ability to interpret and integrate external information.
In Section 2 we provide background on the semantic formalism of our choice, earlier approach to semantic parsing, and the challenge of an important task, word sense disambiguation. Then we propose Retrieval-Augmented Semantic Parsing (RASP) in Section 3, a technique that integrates a retrieval mechanism into parsing. RASP leverages external lexical knowledge in the input, enabling the model to dynamically access and interpret relevant concept information. By incorporating this retrieval step (Section 4), we relax the reliance on lemma-based mappings and allow the model to adapt more naturally to unseen words or senses. Our results show that this approach nearly doubles the performance on predicting OOD concepts compared to previous methods, demonstrating a substantial advancement in handling challenging open-domain data (Section 5).
@@ 2.1 Discourse Representation Structure
Figure 1: Three formats of Discourse Representation Structure (DRS) for ”Mary didn’t commit a crime.”: the box notation, a directed acyclic graph, and the sequence notation.
@@ 2.2 Semantic Parsing
Early approaches to semantic parsing primarily relied on rule-based systems Woods (1973); Hendrix et al. (1977); Templeton and Burger (1983). The advent of neural methodologies, coupled with the availability of large semantically annotated datasets Banarescu et al. (2013); Bos et al. (2017); Abzianidze et al. (2017), marked a significant shift in semantic parsing techniques Barzdins and Gosko (2016); van Noord and Bos (2017); Bevilacqua et al. (2021a). The introduction of pre-trained language models within the sequence-to-sequence framework further improved parsing performance van Noord et al. (2018, 2020); Ozaki et al. (2020); Samuel and Straka (2020); Shou and Lin (2021); Bevilacqua et al. (2021a); Zhou et al. (2021); Mart´ınez Lorenzo et al. (2022); Zhang et al. (2024); Liu (2024a, b); Yang et al. (2024); Amin et al. (2025). Furthermore, several studies introduced more pre-training tasks specifically designed for semantic parsing Bai et al. (2022); Wang et al. (2023a). With the rise of large language models, there has been considerable discussion about leveraging these models for semantic parsing, achieving notable results through techniques like prompting and chain-of-thought reasoning Roy et al. (2022); Ettinger et al. (2023); Jin et al. (2024). However, there is currently no work that leverages the knowledge and understanding capabilities of large language models to address the generalization problem in semantic parsing.
@@ 2.3 Word Sense Disambiguation
The generalization problem introduced in the previous section can also be understood as word sense disambiguation (WSD) for out-of-distribution concepts, within the context of semantic parsing. For instance, consider the sentence ”She had £10,000 in the bank”, with the target word ”bank”. In traditional WSD tasks, a predefined inventory of possible senses (e.g., 1. sloping land; 2. financial institution; 3. a long ridge or pile; 4. …) would be provided, and the WSD model would classify the word according to one of these senses Navigli (2009); Bevilacqua et al. (2021b).
In semantic parsing, WSD can be seen as a subtask Zhang et al. (2025), but it is more challenging because the parsing model must generate the correct sense directly without access to an explicitly provided sense inventory. However, traditional knowledge-based WSD offers a potential solution that inspires our approach: by retrieving and presenting all possible concepts as alternatives, we can explicitly provide external information to the model, thereby enhancing its generalization capability. As a consequence, this requires the model to be able to process long contexts, making the LLMs become the preferences, in particular retrieval augmented generation.
@@ 3 Retrieval-Augmented Semantic Parsing
Gold DRS female.n.02 Name ”Mary” time.n.08 TPR now birdwatch.v.01 Agent -2 Time -1 ELABORATION ¡1 female.n.02 ANA -3 see.v.01 Experiencer -1 Time +1 Stimulus +3 time.n.08 TPR now harrier.n.03 golden_eagle.n.01 entity.n.01 Sub -2 Sub -1 Sub +1 hobby.n.03
@@
Different from the Dense Passage Retrieval Karpukhin et al. (2020) method, which is commonly employed in question-answering tasks, our retrieval process is designed to be more straightforward and tailored to the needs of semantic parsing. The process begins with tokenizing and lemmatizing [^1] https://www.nltk.org/api/nltk.stem.wordnet.html the source text. Following these, we perform a search for relevant concept synsets in an external knowledge base, specifically WordNet. For example, in the sentence ”Mary went for birdwatching. She saw a harrier, a golden eagle, and a hobby”, the retrieval process would identify multiple synsets for ”go”, ”birdwatch”, ”see”, ”harrier”, ”golden eagle” and ”hobby”, as illustrated in Table 1. Additionally, to ensure comprehensive coverage of multi-word expressions, which are critical in capturing the correct semantic meaning, we employ a hierarchical n-gram search strategy. This strategy involves sequential searches using 4-gram, 3-gram, 2-gram, and 1-gram patterns, thereby ensuring that no multi-word expressions (such as ”golden eagle”) are overlooked.
@@
By incorporating retrieved concepts, RASP goes beyond relying solely on the input sequence and training data, adding additional context to guide generation. For example, when handling words with multiple meanings, like ”hobby,” retrieved synsets help the model select the correct interpretation based on glosses and context. This integration sharpens the model’s focus on relevant concepts, reducing the likelihood of generating incorrect or overly broad outputs, particularly for out-of-distribution concepts.
@@ 4.1 Datasets
We conduct our experiments on the Parallel Meaning Bank (PMB, version 5.1.0) [^3] https://pmb.let.rug.nl/releases Abzianidze et al. (2017); Zhang et al. (2024). We first use the gold-standard English data of the PMB to evaluate the large language models and their retrieval-augmented version under in-distribution conditions.
To further assess the models’ ability to handle out-of-distribution (OOD) concepts, we adopt the challenge set proposed by Zhang et al. (2025), which is also derived from the PMB. Neural semantic parsers often default to the first sense of unknown concepts–an approach that can lead to ”lucky guesses” without truly understanding new words. The challenge set, consisting of 500 sentences, is deliberately designed to eliminate this shortcut. Each sentence includes at least one concept that does not appear in the training data and does not correspond to the first sense in the ontology. In total, the challenge set contains 410 unknown nouns, 128 verbs, and 65 modifiers (adjectives and adverbs). By evaluating on this set, we measure the true generalization capability of the models, testing whether they can correctly interpret novel concepts rather than relying on heuristic assigning.
@@ 4.2 Experiment Settings
It is crucial to note that large language models, when used in zero-shot or few-shot scenarios, tend to perform poorly on the highly complex graph structures inherent in formal meaning representations such as DRS. Prior work Ettinger et al. (2023); Zhang et al. (2025) demonstrates that without fine-tuning, LLMs struggle to match the performance of models specifically optimized for these tasks. Therefore, in our experiments, we fine-tune all large language models.
@@
Due to computational constraints, we select open-sourced LLMs with model sizes under 10B parameters, including phi3-4B, Mistral-7B, LLaMa3-3B, LLaMa3-8B, Gemma2-2B, Gemma2-9B, Qwen2.5-3B, and Qwen2.5-7B. These models strike a balance between state-of-the-art language understanding and manageable resource requirements. For fine-tuning, we employ Low-Rank Adaptation (Hu et al., 2021, LoRA), a parameter-efficient technique that introduces trainable low-rank matrices into the model’s layers, greatly reducing computational overhead.
We compare our results against several strong baselines, including BART, T5, byT5, TAX-parser Zhang et al. (2024), and AMS-Parser Yang et al. (2024), all of which were previously fine-tuned on PMB data. We exclude work conducted on earlier versions of PMB or using silver data. Additionally, we do not apply retrieval augmentation to these baseline models due to input length constraints, which limit their ability to incorporate external lexical sources efficiently.
@@
Table 3: Performance of baseline models, large language models (Normal) and their retrieval-augmented variants (RASP) on standard test, with percentage changes in parentheses. Size is the number of model’s parameters (B: billion). IFR is Ill-Formed Rate and the number of ill-formed prediction are in parentheses. Note: AMS-Parser Yang et al. (2024) performs well for IFR for it is a compositional neuro-symbolic system. TAX-Parser Zhang et al. (2025) is a neuro-symbolic system, trained with a novel encoded meaning representation.
@@ 4.3 Evaluation Metrics
We used SMATCH and its variants to evaluate the performance of the models. SMATCH Cai and Knight (2013), referred to as Hard-SMatch, strictly matches concepts, where any discrepancy results in a non-match. In contrast, its variant, Soft-SMatch Opitz et al. (2020), considers concept similarity when matching. Instead of adopting the approach of using word-embedding similarity, we applied the Wu-Palmer similarity Wu and Palmer (1994), as introduced by Zhang et al. (2024). Wu-Palmer similarity provides a precise measure of semantic similarity between concepts based on their positions within the WordNet taxonomy. Unlike embedding-based methods, it does not rely on external training and easily adapts to changes in WordNet’s structure or content. The calculation is:
@@ 5.2 Performance on the Challenge Set
Among the baselines, TAX-Parser Zhang et al. (2025) stands out, achieving an overall similarity score of 39.68. However, some Normal (non-RASP) large language models already exceed this performance on the challenge set. For example, Gemma2-9B (Normal) obtains an overall score of 42.54, indicating that LLMs can yield improvements, even without retrieval augmentation. When retrieval augmentation (RASP) is introduced, these large language models show substantial additional gains. For example, Gemma2-9B (RASP) achieves an overall similarity score of 70.41, compared to the best baseline’s 39.68—an increase of over 30 absolute points. These gains are particularly remarkable for noun concepts, with relative improvements of approximately 70% to 95%. Verbs show increases between about 25% and 77%, and modifiers improve by roughly 26% to 43%.
These results directly support our second research question regarding improving out-of-distribution generalization. While model scaling alone can yield moderate improvements, the integration of external lexical knowledge through retrieval allows LLMs to select more accurate concepts in OOD scenarios. In effect, RASP helps the models ”look up” relevant information, enhancing their concept selection and producing more semantically appropriate results. In this case, retrieval-augmented LLMs not only outperform strong baselines like TAX-Parser but also set the state-of-the-art for OOD semantic parsing performance.
@@ 5.3 Error Analysis on the Challenge Set
We picked 22 instances, comprising 11 completely perfect predictions (WuP=1.00) and 11 imperfect predictions (WuP¡1.00) made by RASP, as presented in Table 5. With respect to the perfect predictions, it is evident that the retrieval significantly enhance the model’s ability to interpret most out-of-distribution concepts. For instance, in the text about birdwatching, the word ”hobby” clearly refers to a species of bird. The model without RAG defaults to the most frequent sense number, predicts hobby.n.01 (an auxiliary activity). In contrast, retrieval provides the glosses of each sense related to the noun ”hobby” and leads the model to pick hobby.n.03 (a falcon), by explicit lexical connections between ”falcon” in the gloss of hobby.n.03 and the context provided by ”birdwatching”.
@@
The verb ”alarm” in Table 5 is an instance of the similarity problem. The challenge arises because some of its senses have similar glosses, such as alarm.v.01 (fill with apprehension or alarm) and alarm.v.02 (warn or arouse to a sense of danger). Similar issues occur with the verbs ”wax”, ”order”, ”observe” and ”hobble”. Although glosses were carefully crafted by lexicographers, they don’t always show a clear difference in meaning Mihalcea and Moldovan (2001); Navigli (2006).
In cases of insufficient textual context, such as with the noun ”kite” in the sentence ”We saw a kite gliding in the sky”, the sense annotators chose kite.n.04 (a bird of prey). However, kite.n.03 (a plaything) could perhaps also be appropriate given the limited context provided by this sentence. Similar issues can be raised in the sentences with the noun ”tripper” and the verb ”feathering”.
The third cause can be attributed to the model’s linguistic coverage. A case in point is ”pen”: the meanings of pen.n.01 (a writing implement) and pen.n.05 (a female swan) are quite different, but the latter is the correct one in the text ”Jane saw two swans. The elegant pen glided gracefully across the tranquil lake”. However, the model fails to distinguish them, likely because ”pen” is rarely used to refer to ”swan” in available corpora. As a result, the models may not have encountered this sense during training, making it challenging for them to predict a meaning they have not been exposed to. In sum, while retrieval drastically improves concept prediction, there are still some difficulties that can pose challenges for the models.